OpenStudy (anonymous):

PLEASEE HELP :) The distance between A (-5, 3) and B(5, 3) is. A. 3 B. 6 C. 10

OpenStudy (anonymous):

C

OpenStudy (anonymous):

distance between two pints (x1,y1) and (x1,y2) \[\sqrt{(x2-x1)^{2} +(y2-y1)^{2}}\]

OpenStudy (anonymous):

Thank you! do you think you could help me with one more problem

OpenStudy (anonymous):

If the point (a,3) lies on the graph of the equation 5x + y = 8, then a= A. 1 B. -1 C. -7

OpenStudy (anonymous):

as (a,3) lies on the graph of the equation , so u need to plug in x=a and y=3 in 5x+y = 8

OpenStudy (anonymous):

So it would be A?